# Optimal Convergence Rate for Mirror Descent Methods with special Time-Varying Step Sizes Rules

### v5

@@ 1. Introduction
The paper consists of an introduction and 5 main sections. In Sect. 2 we mentioned the basic facts and tools and mirror descent basics. Sect. 3 devoted to the proof of the optimal convergence rate of the mirror descent algorithm with time-varying step size rules for solving the classical constrained non-smooth optimization problem. In Sect. 4 the same analysis, that in Sect. 3, performed for more general optimization problems, namely the constrained composite optimization problem. Section 5 is devoted to the problem with functional constraints, which is a more general class of non-smooth optimization problems. In Sect. 6 we present some numerical experiments which demonstrate the effectiveness of the proposed algorithms and analysis.
@@ 2. Fundamentals
(2.3) $\left|\langle a,b\rangle\right|\leq\frac{\|a\|^{2}}{2\lambda}+\frac{\lambda\|b\|_{*}}{2},\quad\forall\lambda>0.$
@@ 3. Mirror descent method for constrained problem
As a special case of Theorem 3.1, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate $O\left(\frac{\log(N)}{\sqrt{N}}\right)$ of Algorithm 1, see [4, 12, 19, 24], with the following (adaptive and non-adaptive) time-varying step size rules
(3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\quad\text{or}\quad\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\quad k=1,2,\ldots,N.$
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=-1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
@@
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, $V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.9) we find
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, $V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.9) we find
$f(\tilde{x})-f(x^{*})$ $\leq\frac{\theta+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}=\frac{M_{f}}{\sqrt{2\sigma}}\frac{\theta+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\theta+1+\log(N)}{\sqrt{N}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}}\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}x^{k}$. ∎
Also, from Theorem 3.1, with a special value of the parameter $m$, we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 1, with the time-varying step sizes given in (3.7).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=0$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
@@
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.11) we find
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.11) we find
$f(\overline{x})-f(x^{*})$ $\leq\frac{1}{N}\left(\frac{\theta\sqrt{N}\|\nabla f(x^{N})\|_{*}}{\sqrt{2\sigma}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{\sqrt{k}}\|\nabla f(x^{k})\|_{*}\right)$
$\leq\frac{1}{N}\left(\frac{\theta M_{f}\sqrt{N}}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}\theta+2\sqrt{N}\right)$
$=\frac{M_{f}(2+\theta)}{\sqrt{2\sigma}}\cdot\frac{1}{\sqrt{N}}.$
Now, from Theorem 3.1, with any fixed $m\geq 1$, we can obtain the optimal convergence rate of Algorithm 1 $O\left(\frac{1}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.7).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with any $m\geq 1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
@@
Let us see the non-adaptive rule (in a similar way we can consider the adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, $V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.4) we find
@@ 4. Mirror descent method for constrained composite problem
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with a positive non-increasing sequence of step sizes $\{\gamma_{k}\}_{k\geq 1}$, for any fixed $m\geq-1$, it satisfies the following inequality
@@
Because of the sequence $\{\gamma_{k}\}_{k\geq 1}$ non-increasing, we have $\frac{1}{\gamma_{k-1}^{m}}\geq\frac{1}{\gamma_{k}^{m}}$, and then we conclude
@@
As a special case of Theorem 4.3, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of the mirror-C descent method (Algorithm 2), namely $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.7) (see [4, 12, 19, 24]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with $m=-1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.13), we find
$F(\tilde{x})-F(x^{*})$ $\leq\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+\theta+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}}$
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+1+\theta+\log(N)}{\sqrt{N}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}}\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}x^{k}$. ∎
Also, from Theorem 4.3, with a special value of the parameter $m$, we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 2, with the time-varying step sizes given in (3.7).
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with $m=0$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
@@
Case 1 (non-adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.14) we find
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.14) we find
$F(\overline{x})-F(x^{*})\leq\frac{1}{N}\left(\frac{\theta M_{f}\sqrt{N}}{\sqrt{2\sigma}}+\frac{M_{f}}{\sqrt{2\sigma}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)\leq\frac{M_{f}(2+\theta)}{\sqrt{2\sigma}}\cdot\frac{1}{\sqrt{N}}.$
Remark 4.6.
In a similar way as in Corollary 3.4, for any $m\geq 1$ in (4.3) with the assumption that $h(x^{1})=0$ we get the same result (3.12), and the convergence rate of Algorithm 2 with the time-varying step sizes given in (3.7), will be optimal. In comparison with the sub-optimal convergence rate (4.12), when $m\geq 1$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points that generated by Algorithm 2.
@@ 5. Mirror descent method for problems with functional constraints
where $f$ and $g_{i}$ (for all $i=1,\ldots,p$) are convex and non-smooth functions given on a compact convex set $Q$.
@@
The output point (output of Algorithm 3) will be selected among the points $x^{k}$ for which $g(x^{k})\leq\varepsilon$. Therefore, we will call step $k$ productive if $g(x^{k})\leq\varepsilon$. If the reverse inequality $g(x^{k})>\varepsilon$ holds then step $k$ will be called non-productive. Let $I$ and $J$ denote the set of productive and non-productive steps, respectively. $|I|$ and $|J|$ denote the number of productive and non-productive steps, respectively. Let us also set $\gamma_{k}:=\gamma_{k}^{f}$ if $k\in I$, $\gamma_{k}:=\gamma_{k}^{g}$ if $k\in J$, and $M=\max\{M_{f},M_{g}\}$.
@@
where $\hat{x}=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$, and $x^{*}$ is an optimal solution of (5.2).
@@
Therefore, by setting $\hat{x}:=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$, we get
@@
Remark 5.3.
Let $x^{1}\in Q$ be an initial point of Algorithm 3, such that $V_{\psi}(x^{*},x^{1})\leq\theta$ for some $\theta>0$. From inequality (5.7) (which is equivalent to the inequality (5.3)), with $\gamma_{q}=\gamma_{q}^{f}$ if $q\in I$ and $\gamma_{q}=\gamma_{q}^{g}$ if $q\in J$, with
@@
$\quad\qquad\leq\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}-\Bigg{(}\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}-\frac{\theta}{\gamma_{k}^{m+1}}-\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}$
@@
(5.11) $\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}\geq\frac{\theta}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g}\right)^{m-1}},$
@@
(5.12) $\gamma_{k}=\begin{cases}\gamma_{k}^{f}:=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\quad\text{or}\quad\gamma_{k}^{f}:=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}};&\text{if}\;k\in I,\\ \gamma_{k}^{g}:=\frac{\sqrt{2\sigma}}{M_{g}\sqrt{k}},\quad\text{or}\quad\gamma_{k}^{g}:=\frac{\sqrt{2\sigma}}{\|\nabla g(x^{k})\|_{*}\sqrt{k}};&\text{if}\;k\in J.\end{cases}$
Let us assume that $V_{\psi}(x^{*},x^{1})\leq\theta,$ for some $\theta>0$, $M:=\max\{M_{f},M_{g}\}$. By using the non-adaptive rules from (5.12) (we also can conclude the same results if we take the adaptive step size rules), and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}\leq M$ and $\|\nabla g(x^{k})\|_{*}\leq M_{g}\leq M$, then for any $m\geq 1$, from Theorem 5.1 we have
$f(\hat{x})-f(x^{*})$ $<\frac{\left(\sqrt{2\sigma}\right)^{m}}{M^{m}\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\Bigg{(}\frac{\theta M^{m+1}\left(\sqrt{N}\right)^{m+1}}{\left(\sqrt{2\sigma}\right)^{m+1}}+$
@@
$=\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta\left(\sqrt{N}\right)^{m+1}+\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m-1}\right)$
$\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta\left(\sqrt{N}\right)^{m+1}+N\left(\sqrt{N}\right)^{m-1}\right)$
$=\frac{M(1+\theta)\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}.$
Now, by setting $\frac{M(1+\theta)\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon$ and since $|I|\leq N$, we get
$\frac{M(1+\theta)\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}N\left(\sqrt{N}\right)^{m}}\leq\frac{M(1+\theta)\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon.$
@@
$\frac{M(1+\theta)}{\sqrt{N}\sqrt{2\sigma}}\leq\varepsilon\quad\Longrightarrow\quad N\geq\frac{M^{2}(1+\theta)^{2}}{2\sigma\varepsilon^{2}}.$
@@
Let $f$ and $g$ are Lipschitz convex functions with constants $M_{f}>0$ and $M_{g}>0$, respectively. Assume that $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g(x)\|_{*}\leq M_{g}$, for all $x\in Q$, and $V_{\psi}(x^{*},x^{1})\leq\theta,$ for some $\theta>0$, and let $\varepsilon>0,M:=\max\{M_{f},M_{g}\}$. Then after
(5.13) $N=\left\lceil\frac{M^{2}(1+\theta)^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right)$
@@
where $\hat{x}=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$.

### v6

@@ 1. Introduction
The paper consists of an introduction and 5 main sections. In Sect. 2 we mentioned the basic facts and tools and mirror descent basics. Sect. 3 devoted to the proof of the optimal convergence rate of the mirror descent algorithm with time-varying step size rules for solving the classical constrained non-smooth optimization problem. In Sect. 4, the same analysis that in Sect. 3, performed for more general optimization problems, namely the constrained composite optimization problem. Section 5 is devoted to the problem with functional constraints, which is a more general class of non-smooth optimization problems. In Sect. 6, we present some numerical experiments which demonstrate the effectiveness of the proposed algorithms and analysis.
@@ 2. Fundamentals
(2.3) $\langle a,b\rangle\leq\frac{\|a\|^{2}}{2\lambda}+\frac{\lambda\|b\|_{*}^{2}}{2},\quad\forall\lambda>0.$
@@ 3. Mirror descent method for constrained problem
As a special case of Theorem 3.1, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate $O\left(\frac{\log(N)}{\sqrt{N}}\right)$ of Algorithm 1, see [4, 12, 19, 24], with the following time-varying step size rule
(3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}\quad k=1,2,\ldots,N.$
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=-1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step size given in (3.7), it satisfies the following sub-optimal convergence rate
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, $V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.9) we find
@@
Also, from Theorem 3.1, with a special value of the parameter $m$, we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 1, with the time-varying step size given in (3.7).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=0$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step size given in (3.7), it satisfies the following optimal convergence rate
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.11) we find
@@
Now, from Theorem 3.1, with any fixed $m\geq 1$, we can obtain the optimal convergence rate of Algorithm 1 $O\left(\frac{1}{\sqrt{N}}\right)$, with the time-varying step size given in (3.7).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with any $m\geq 1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step size given in (3.7), it satisfies the following optimal convergence rate
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, $V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (3.4) we find
@@ 4. Mirror descent method for constrained composite problem
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with a positive non-increasing sequence of step sizes $\{\gamma_{k}\}_{k\geq 1}$, for any fixed $-1\leq m\leq 0$, it satisfies the following inequality
@@
Because of the sequence $\{\gamma_{k}\}_{k\geq 1}$ non-increasing and $-1\leq m\leq 0$, we have $\frac{1}{\gamma_{k-1}^{m}}\geq\frac{1}{\gamma_{k}^{m}}$, and then we conclude
@@
As a special case of Theorem 4.3, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of the mirror-C descent method (Algorithm 2), namely $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, with the time-varying step size given in (3.7) (see [4, 12, 19, 24]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with $m=-1$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step size given in (3.7), it satisfies the following sub-optimal convergence rate
@@
Also, from Theorem 4.3, with a special value of the parameter $m$, we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 2, with the time-varying step size given in (3.7).
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with $m=0$, $V_{\psi}(x^{*},x^{1})\leq\theta$, for some $\theta>0$, and with the time-varying step size given in (3.7), it satisfies the following optimal convergence rate
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f},V_{\psi}(x^{*},x^{1})\leq\theta$, then by substitution in (4.14) we find
@@ 5. Mirror descent method for problems with functional constraints
where $f$ and $g_{i}$ (for all $i=1,\ldots,p$) are convex and non-smooth functions given on a compact convex set $Q$, this means that they are Lipschitz functions, i.e., there is $M_{f}>0$ and $M_{g_{i}}>0\;(i=1,2,\ldots,p)$, such that $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g_{i}(x)\|_{*}\leq M_{g_{i}}(i=1,2,\ldots,p)$.
@@
The output point (output of Algorithm 3) will be selected among the points $x^{k}$ for which $g(x^{k})\leq\varepsilon$. Therefore, we will call step $k$ productive if $g(x^{k})\leq\varepsilon$. If the reverse inequality $g(x^{k})>\varepsilon$ holds then step $k$ will be called non-productive. Let $I$ and $J$ denote the set of productive and non-productive steps, respectively. $|I|$ and $|J|$ denote the number of productive and non-productive steps, respectively. Let us also set $\gamma_{k}:=\gamma_{k}^{f}$ if $k\in I$, $\gamma_{k}:=\gamma_{k}^{g}$ if $k\in J$, and $M=\max\{M_{f},M_{g}\}$, where $M_{g}=\max_{1\leq i\leq p}\{M_{g_{i}}\}$.
@@
where $\hat{x}=\frac{1}{\sum\limits_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$, and $x^{*}$ is an optimal solution of (5.2).
@@
Therefore, by setting $\hat{x}:=\frac{1}{\sum\limits_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$, we get
@@
Remark 5.3 (Stopping criterion for Algorithm 3).
Let $x^{1}\in Q$ be an initial point of Algorithm 3, such that $V_{\psi}(x^{*},x^{1})\leq\theta_{1}$ for some $\theta_{1}>0$. From inequality (5.7) (which is equivalent to the inequality (5.3)), with $\gamma_{q}=\gamma_{q}^{f}$ if $q\in I$ and $\gamma_{q}=\gamma_{q}^{g}$ if $q\in J$, with
@@
$\quad\qquad\leq\varepsilon\sum_{i\in I}\left(\gamma_{i}^{f}\right)^{-m}-\Bigg{(}\varepsilon\sum_{i=1}^{k}\left(\gamma_{i}\right)^{-m}-\frac{\theta_{1}}{\gamma_{k}^{m+1}}-\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}$
@@
(5.11) $\varepsilon\sum_{i=1}^{k}\frac{1}{\gamma_{i}^{m}}\geq\frac{\theta_{1}}{\gamma_{k}^{m+1}}+\frac{1}{2\sigma}\sum_{i\in I}\frac{\|\nabla f(x^{i})\|_{*}^{2}}{\left(\gamma_{i}^{f}\right)^{m-1}}+\frac{1}{2\sigma}\sum_{j\in J}\frac{\|\nabla g(x^{j})\|_{*}^{2}}{\left(\gamma_{j}^{g}\right)^{m-1}},$
@@
(5.12) $\gamma_{k}=\begin{cases}\gamma_{k}^{f}:=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\quad\text{if}\;k\in I,\\ \gamma_{k}^{g}:=\frac{\sqrt{2\sigma}}{M_{g}\sqrt{k}},\quad\text{if}\;k\in J.\end{cases}$
Let us assume that $V_{\psi}(x^{*},x^{1})\leq\theta_{1},$ for some $\theta_{1}>0$, $M:=\max\{M_{f},M_{g}\}$. By using (5.12), and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}\leq M$ and $\|\nabla g(x^{k})\|_{*}\leq M_{g}\leq M$, then for any $m\geq 1$, from Theorem 5.1 we have
$f(\hat{x})-f(x^{*})$ $<\frac{\left(\sqrt{2\sigma}\right)^{m}}{M^{m}\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\Bigg{(}\frac{\theta_{1}M^{m+1}\left(\sqrt{N}\right)^{m+1}}{\left(\sqrt{2\sigma}\right)^{m+1}}+$
@@
$=\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta_{1}\left(\sqrt{N}\right)^{m+1}+\sum_{k=1}^{N}\left(\sqrt{k}\right)^{m-1}\right)$
$\leq\frac{M}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\cdot\left(\theta_{1}\left(\sqrt{N}\right)^{m+1}+N\left(\sqrt{N}\right)^{m-1}\right)$
$=\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}.$
Now, by setting $\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon$ and since $|I|\leq N$, we get
$\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}N\left(\sqrt{N}\right)^{m}}\leq\frac{M(1+\theta_{1})\left(\sqrt{N}\right)^{m+1}}{\sqrt{2\sigma}}\cdot\frac{1}{\sum_{k\in I}\left(\sqrt{k}\right)^{m}}\leq\varepsilon.$
@@
$\frac{M(1+\theta_{1})}{\sqrt{N}\sqrt{2\sigma}}\leq\varepsilon\quad\Longrightarrow\quad N\geq\frac{M^{2}(1+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}.$
@@
Let $f$ and $g$ are Lipschitz convex functions with constants $M_{f}>0$ and $M_{g}>0$, respectively. Assume that $\|\nabla f(x)\|_{*}\leq M_{f}$ and $\|\nabla g(x)\|_{*}\leq M_{g}$, for all $x\in Q$, and $V_{\psi}(x^{*},x^{1})\leq\theta_{1},$ for some $\theta_{1}>0$, and let $\varepsilon>0,M:=\max\{M_{f},M_{g}\}$. Then after
(5.13) $N=\left\lceil\frac{M^{2}(1+\theta_{1})^{2}}{2\sigma\varepsilon^{2}}\right\rceil=O\left(\frac{1}{\varepsilon^{2}}\right)$
@@
where $\hat{x}=\frac{1}{\sum_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}}\sum\limits_{k\in I}\left(\gamma_{k}^{f}\right)^{-m}x^{k}$.